# A Theoretical Framework for Grokking: Interpolation followed by Riemannian Norm Minimisation

### v1

@@ A.4 Proof of Proposition 3
By definition [see e.g., Bergmann and Herzog, 2019], the KKT points of Equation 7 are the points $w^{\star}\in\mathcal{M}$ satisfying
@@
Since $\mathrm{grad}_{\mathcal{M}}\ell_{2}=P_{\mathrm{Ker}(\nabla^{2}F(w^{\star}))}$, KKT points of Equation 7 are the points $w^{\star}\in\mathcal{M}$ satisfying
@@
Thanks to Lemma 7, the trajectories $(w^{\lambda}(t))_{t\geq 0}$ are all bounded and $w^{\lambda}_{\infty}\coloneqq\lim_{t\to\infty}w^{\lambda}(t)$ exists for any $\lambda>0$. In particular, this limit is a stationary point of the regularised loss $F_{\lambda}$, i.e.,
@@
Let $(\lambda_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathbb{R}_{+}^{*}$ such that $\lambda_{k}\underset{k\to\infty}{\longrightarrow}0$. Let $w^{\star}$ be an adherence point of the sequence $(w^{\lambda_{k}}_{\infty})_{k}$. Thanks to Lemma 3, $w^{\star}\in\mathcal{M}$. Moreover, the equality $w^{\lambda_{k}}_{\infty}=-\frac{1}{\lambda_{k}}\nabla F(w^{\lambda_{k}}_{\infty})$ first implies that $\|\nabla F(w^{\lambda_{k}}_{\infty})\|=\mathcal{O}\left(\lambda_{k}\right)$. Rebjock and Boumal [2024, Proposition 2.8] then also implies that
@@
Moreover, noting $w_{k}=\mathrm{arg\ min\ }_{w\in\mathcal{M}}\|w^{\lambda_{k}}_{\infty}-w\|$ the projection of $w^{\lambda_{k}}_{\infty}$ onto $\mathcal{M}$, a Taylor expansion yields
$\frac{1}{\lambda}\nabla F(w^{\lambda_{k}}_{\infty})$ $=\frac{1}{\lambda}\nabla F(w_{k})+\frac{1}{\lambda}\nabla^{2}F(w_{k})^{\top}(w^{\lambda_{k}}_{\infty}-w_{k})+o(\frac{\|w^{\lambda_{k}}-w_{k}\|}{\lambda_{k}})$
$=\frac{1}{\lambda}\nabla^{2}F(w_{k})^{\top}(w^{\lambda_{k}}_{\infty}-w_{k})+o(\frac{d(w^{\lambda_{k}}_{\infty},\mathcal{M})}{\lambda_{k}})$
$=\frac{1}{\lambda}\nabla^{2}F(w_{k})^{\top}(w^{\lambda_{k}}_{\infty}-w_{k})+o(1).$
The equality $w^{\lambda_{k}}_{\infty}=-\frac{1}{\lambda_{k}}\nabla F(w^{\lambda_{k}}_{\infty})$ then implies that
$-\lim_{k\to\infty}\nabla^{2}F(w_{k})^{\top}\frac{w^{\lambda_{k}}_{\infty}-w_{k}}{\lambda_{k}}=w^{\star}.$ (13)
Let $u_{k}=P_{\mathrm{Ker}(\nabla^{2}F(w_{k}))}\frac{w^{\lambda_{k}}_{\infty}-w_{k}}{\lambda_{k}}$. Note that
$\nabla^{2}F(w_{k})^{\top}\frac{w^{\lambda_{k}}_{\infty}-w_{k}}{\lambda_{k}}=\nabla^{2}F(w_{k})^{\top}u_{k}$
$\text{and}\quad\|\nabla^{2}F(w_{k})^{\top}u_{k}\|\geq\eta\|u_{k}\|,$
thanks to Assumption 2. In consequence, $(u_{k})_{k}$ is bounded. In particular, it admits an adherence point $u_{\infty}\in\mathbb{R}^{d}$.
Since $w_{k}\underset{k\to\infty}{\rightarrow}w^{\star}$ and $\nabla^{2}F$ is continuous, $\|\nabla^{2}F(w_{k})-\nabla^{2}F(w^{\star})\|\underset{k\to\infty}{\rightarrow}0$. So that Equation 13 becomes
$-\nabla^{2}F(w^{\star})^{\top}u_{\infty}=w^{\star}.$
In particular, it yields that $w^{\star}\in\mathrm{Im}(\nabla^{2}F(w^{\star}))$. By symmetry of the Hessian, $\mathrm{Im}(\nabla^{2}F(w^{\star}))=\mathrm{Ker}(\nabla^{2}F(w^{\star}))^{\perp}$ so that the adherence point $w^{\star}$ of $(w^{\lambda_{k}}_{\infty})_{k}$ satisfies the KKT conditions of Equation 7, which are equivalent to ∎
A.5 Proof of Proposition 4
@@
For this proof, denote $w^{\star}=\lim_{t\to\infty}\tilde{w}^{\circ}(t)$ and $F^{\star}=F(w^{\star})$. $w^{\star}$ is a strict local minimum of the Euclidean norm on $\mathcal{M}$. Moreover using the Morse Bott property [Assumption 2 and see Rebjock and Boumal, 2024], we can consider an arbitrarily small $\delta>0$ such that the following conditions simultaneously hold in $B(0,\delta)$ for some $\beta>0$:
$\forall w\in\mathcal{M}\cap B(0,2\delta),w\neq w^{\star}\implies\|w^{\star}\|^{2}<\|w\|^{2},$
$\forall w\in B(0,\delta),F(w)-F^{\star}\geq\frac{\eta}{4}d(w,\mathcal{M})^{2},$ (14)
$\forall w\in B(0,\delta),\beta(F(w)-F^{\star})\geq\|\nabla F(w)\|_{2}^{2}\geq\eta(F(w)-F^{\star}).$
First observe that the strict minimality assumption implies, through Lemma 8, that there exists $\varepsilon_{0}>0$ (independent of $\lambda$) such that for a small enough $\lambda>0$
@@
From there, we define for this proof $T_{\lambda}=\inf\{t\geq t_{0}\mid\tilde{w}^{\lambda}(t)\not\in B(w^{\star},\delta^{\prime})\}$. Similarly to the proof of Lemma 7, we have for any $t\in(\frac{t_{0}}{\lambda},\frac{T_{\lambda}}{\lambda})$:
@@
$F_{\lambda}(w^{\lambda}(t))\leq\varepsilon e^{-\eta(t-\frac{t_{0}}{\lambda})}+\lambda\left(\frac{R^{2}}{2}+\frac{R_{1}}{\eta}\right).$
In particular, if we define $t^{\prime}=\min(\frac{t_{0}}{\lambda}-\frac{\ln(\lambda)}{\eta},\frac{T_{\lambda}}{\lambda})$, we have similarly to the proof of Lemma 7 that for any $t\in[\frac{t_{0}}{\lambda},t^{\prime}]$:
@@
for some constant $C$ independent of $\varepsilon$ and $\lambda$. In particular, we can choose $\varepsilon$ and $\lambda$ small enough so that this quantity is smaller than $\frac{\delta}{2}$. It then implies that $t^{\prime}<\frac{T_{\lambda}}{\lambda}$ and
@@
By continuity, Equation 15 then implies that for any $t\geq t^{\prime}$, $w^{\lambda}(t)\in B(w^{\star},\delta)$.
To summarise, we have shown that for any small enough $\delta>0$, there exists $\lambda^{\star}(\delta)$ and $t^{\prime}(\lambda,\delta)$ such that for any $\lambda\leq\lambda^{\star}(\delta)$, $\lim_{t\to\infty}w^{\lambda}(t)\in B(w^{\star},\delta)$.
This means that $\lim_{\lambda\to 0}\lim_{t\to\infty}w^{\lambda}(t)=w^{\star}$, which proves Proposition 4. ∎
@@ B.1 Proof of Lemma 1
The main point of the proof is to show that $\mathcal{M}\cap B$ is geometrically stable in the sense of Falconer [1983], i.e., that there exists a neighbourhood (in $B$) $U$ of $\mathcal{M}\cap B$, $s>0$ and $k<1$ such that for any $u\in U$ and $t\geq 0$,
$d(\phi(w,t+s),\mathcal{M}\cap B)\leq kd(\phi(w,t),\mathcal{M}\cap B),$
where $d(v,\mathcal{M})=\inf_{x\in\mathcal{M}}\|v-x\|_{2}$. [^3] The definition of Falconer [1983] is stated differently but a consequence of ours, by defining the open set $W=\{\phi(w,t)\mid w\in U,t\geq 0\}$ and $f(w)=\phi(w,s)$.
Let $x\in\mathcal{M}\cap B$. The Morse-Bott property (Assumption 2) implies that there exists a neighbourhood $U(x)\subset B$ of $x$, such that $F$ satisfies the Polyak-Łojasiewicz (PL) inequality with constant $\frac{\eta}{2}$, thanks to the equivalences between both conditions [Rebjock and Boumal, 2024]
@@
In particular, there is some $\delta_{0}(x)>0$ such that $B(x,\delta_{0}(x))\subset U(x)$. Thanks to Rebjock and Boumal [2024, Propositions 2.3 and 2.8, Remark 2.10], we can even choose $\delta_{0}(x)$ small enough so that there are some $\alpha,\beta$ such that
@@
By boundedness of $B$, note that $\alpha$ and $\beta$ can be chosen indepently of $x\in\mathcal{M}\cap B$ here.
Now let $w\in B(x,\delta_{0}(x))$ and define $T(w)=\inf\{t\geq 0\mid\phi(w,t)\not\in B(x,\delta_{0}(x))\}$. Necessarily, $T(w)>0$ and for any $t\in[0,T(w))$, Equation 17 applies to $\phi(w,t)$, so that for any $t\in[0,T(w))$
@@
Moreover for any $t\in[0,T(w))$, Equation 18 also applies, so that
@@
Also, note that $\|\phi(w,t)-x\|\leq\frac{\delta_{0}(x)}{2}$ and $B(x,\delta_{0}(x))\subset B$ implies that $d(\phi(w,t),\mathcal{M}\cap B)=d(\phi(w,t),\mathcal{M})$. From there, Equation 19 implies for any $t\geq 0$ and $w\in B(x,\delta(x))$:
@@
In particular, for a sufficiently large $t$, $d(\phi(w,t),\mathcal{M}\cap B)\leq kd(w,\mathcal{M}\cap B)$ for some $k<1$. Then taking $U(B)=\bigcup_{x\in\mathcal{M}\cap B}B(x,\delta(x))$, $\mathcal{M}\cap B$ is geometrically stable for $f:w\mapsto\phi(w,t)$.
Additionally thanks to our previous remark, $U$ is invariant by $f$. We can now apply Falconer [1983, Lemma 6.2 and Theorem 5.1], which implies that $\Phi$ is $\mathcal{C}^{2}$ on $U(B)$. Taking an increasing sequence of open bounded sets $B_{n}$ covering whole $\mathbb{R}^{d}$, we can then define $U=\bigcup_{n}U(B_{n})$ and conclude that $\Phi$ is $\mathcal{C}^{2}$ on $U$.
@@ B.2 Minimality of $F$ on neighbourhood
Consider Assumptions 1 and 2. Let $U$ be an open neighbourhood of $\mathcal{M}$ such that $\Phi$ is continuous on $U$, then necessarily for any $w\in U\setminus\mathcal{M}$, $F(w)=\sup_{x\in\mathcal{M}}F(x)$.
@@ B.3 Alternative equation for $\tilde{w}^{\circ}$
If Assumption 2 holds, the unique solution of Equation 6 also corresponds to the unique solution of the following equation:
@@
The first one is a consequence of the definition of the Riemannian gradient and the fact that $T_{\mathcal{M}}(w)=\mathrm{Ker}(\nabla^{2}F(w))$ [see e.g., Boumal, 2023, Theorem 3.15 with $\nabla F$ being the local defining function of $\mathcal{M}$]. The second one is given by Li et al. [2021, Lemma 4.3]. ∎

### v2

@@ A.4 Proof of Proposition˜3
By definition [see e.g., bergmann2019intrinsic], the KKT points of Equation˜7 are the points $w^{\star}\in\mathcal{M}$ satisfying
@@
Since $\mathrm{grad}_{\mathcal{M}}\ell_{2}=P_{\mathrm{Ker}(\nabla^{2}F(w^{\star}))}$, KKT points of Equation˜7 are the points $w^{\star}\in\mathcal{M}$ satisfying
@@
Thanks to Lemma˜7, the trajectories $(w^{\lambda}(t))_{t\geq 0}$ are all bounded and $w^{\lambda}_{\infty}\coloneqq\lim_{t\to\infty}w^{\lambda}(t)$ exists for any $\lambda>0$. In particular, this limit is a stationary point of the regularised loss $F_{\lambda}$, i.e.,
@@
Let $(\lambda_{k})_{k\in\mathbb{N}}$ be a sequence in $\mathbb{R}_{+}^{*}$ such that $\lambda_{k}\underset{k\to\infty}{\longrightarrow}0$. Let $w^{\star}$ be a limit point of the sequence $(w^{\lambda_{k}}_{\infty})_{k}$. Thanks to Lemma˜3, $w^{\star}\in\mathcal{M}$. Moreover, the equality $w^{\lambda_{k}}_{\infty}=-\frac{1}{\lambda_{k}}\nabla F(w^{\lambda_{k}}_{\infty})$ first implies that $\|\nabla F(w^{\lambda_{k}}_{\infty})\|=\mathcal{O}\left(\lambda_{k}\right)$. rebjock2024fast then also implies that
@@
Moreover, noting $w_{k}\in\mathrm{arg\ min\ }_{w\in\mathcal{M}}\|w^{\lambda_{k}}_{\infty}-w\|$, a Taylor expansion yields
$\frac{1}{\lambda_{k}}\nabla F(w^{\lambda_{k}}_{\infty})$ $=\frac{1}{\lambda_{k}}\nabla F(w_{k})+\frac{1}{\lambda_{k}}\nabla^{2}F(w_{k})(w^{\lambda_{k}}_{\infty}-w_{k})+o(\frac{\|w^{\lambda_{k}}-w_{k}\|}{\lambda_{k}})$
$=\frac{1}{\lambda_{k}}\nabla^{2}F(w_{k})(w^{\lambda_{k}}_{\infty}-w_{k})+o(\frac{d(w^{\lambda_{k}}_{\infty},\mathcal{M})}{\lambda_{k}})$
$=\frac{1}{\lambda_{k}}\nabla^{2}F(w_{k})(w^{\lambda_{k}}_{\infty}-w_{k})+o(1).$
The equality $w^{\lambda_{k}}_{\infty}=-\frac{1}{\lambda_{k}}\nabla F(w^{\lambda_{k}}_{\infty})$ then implies for the subsequence $k_{n}$ associated to the limit point $w^{\star}$ that
$\lim_{n\to\infty}\nabla^{2}F(w_{k_{n}})\frac{w^{\lambda_{k_{n}}}_{\infty}-w_{k_{n}}}{\lambda_{k_{n}}}=w^{\star}.$ (13)
Let $u_{k}=P_{\mathrm{Ker}(\nabla^{2}F(w_{k}))^{\perp}}\frac{w^{\lambda_{k}}_{\infty}-w_{k}}{\lambda_{k}}$. Note that
$\nabla^{2}F(w_{k})\frac{w^{\lambda_{k}}_{\infty}-w_{k}}{\lambda_{k}}=\nabla^{2}F(w_{k})u_{k}$
$\text{and}\quad\|\nabla^{2}F(w_{k})u_{k}\|\geq\eta\|u_{k}\|,$
thanks to Assumption˜2. In consequence, $(u_{k})_{k}$ is bounded. In particular, it admits an adherence point $u_{\infty}\in\mathbb{R}^{d}$.
Since $w_{k_{n}}\underset{n\to\infty}{\rightarrow}w^{\star}$ and $\nabla^{2}F$ is continuous, $\|\nabla^{2}F(w_{k_{n}})-\nabla^{2}F(w^{\star})\|\underset{n\to\infty}{\rightarrow}0$. So that Equation˜13 becomes
$\nabla^{2}F(w^{\star})u_{\infty}=w^{\star}.$
In particular, it yields that $w^{\star}\in\mathrm{Im}(\nabla^{2}F(w^{\star}))$. By symmetry of the Hessian, $\mathrm{Im}(\nabla^{2}F(w^{\star}))=\mathrm{Ker}(\nabla^{2}F(w^{\star}))^{\perp}$ so that $w^{\star}\in\mathrm{Ker}(\nabla^{2}F(w^{\star}))^{\perp}$, i.e., it satisfies the KKT conditions of Equation˜7. ∎
A.5 Proof of Proposition˜4
@@
For this proof, denote $w^{\star}=\lim_{t\to\infty}\tilde{w}^{\circ}(t)$ and $F^{\star}=F(w^{\star})$. $w^{\star}$ is a strict local minimum of the Euclidean norm on $\mathcal{M}$. Moreover using the Morse Bott property [Assumption˜2 and see rebjock2024fast], we can consider an arbitrarily small $\delta>0$ such that the following conditions simultaneously hold in $B(0,\delta)$ for some $\beta>0$:
$\forall w\in\mathcal{M}\cap B(w^{\star},2\delta),w\neq w^{\star}\implies\|w^{\star}\|^{2}<\|w\|^{2},$
$\forall w\in B(w^{\star},\delta),F(w)-F^{\star}\geq\frac{\eta}{4}d(w,\mathcal{M})^{2},$ (14)
$\forall w\in B(w^{\star},\delta),\beta(F(w)-F^{\star})\geq\|\nabla F(w)\|_{2}^{2}\geq\eta(F(w)-F^{\star}).$
First observe that the strict minimality assumption implies, through Lemma˜8, that there exists $\varepsilon_{0}>0$ (independent of $\lambda$) such that for a small enough $\lambda>0$
@@
From there, we define for this proof $T_{\lambda}=\inf\{t\geq t_{0}\mid\tilde{w}^{\lambda}(t)\not\in B(w^{\star},\delta^{\prime})\}$. Similarly to the proof of Lemma˜7, we have for any $t\in(\frac{t_{0}}{\lambda},\frac{T_{\lambda}}{\lambda})$:
@@
$F_{\lambda}(w^{\lambda}(t))\leq F^{\star}+\varepsilon e^{-\eta(t-\frac{t_{0}}{\lambda})}+\lambda\left(\frac{R^{2}}{2}+\frac{R_{1}}{\eta}\right).$
In particular, if we define $t^{\prime}=\min(\frac{t_{0}}{\lambda}-\frac{\ln(\lambda)}{\eta},\frac{T_{\lambda}}{\lambda})$, we have similarly to the proof of Lemma˜7 that for any $t\in[\frac{t_{0}}{\lambda},t^{\prime}]$:
@@
for some constant $C$ independent of $\varepsilon$ and $\lambda$. In particular, we can choose $\varepsilon$ and $\lambda$ small enough so that this quantity is smaller than $\frac{\delta^{\prime}}{2}$. It then implies that $t^{\prime}<\frac{T_{\lambda}}{\lambda}$ and
@@
By continuity, Equation˜15 then implies that for any $t\geq t^{\prime}$, $w^{\lambda}(t)\in B(w^{\star},\delta)$.
To summarise, we have shown that for any small enough $\delta>0$, there exists $\lambda^{\star}(\delta)$ such that for any $\lambda\leq\lambda^{\star}(\delta)$, $\lim_{t\to\infty}w^{\lambda}(t)\in B(w^{\star},\delta)$.
This means that $\lim_{\lambda\to 0}\lim_{t\to\infty}w^{\lambda}(t)=w^{\star}$, which proves Proposition˜4. ∎
@@ B.1 Proof of Lemma˜1
The main point of the proof is to show that $\mathcal{M}\cap B$ is geometrically stable in the sense of falconer1983differentiation, i.e., that there exists a neighbourhood (in $B$) $U$ of $\mathcal{M}\cap B$, $t>0$ and $k<1$ such that for any $w\in U$,
$d(\phi(w,t),\mathcal{M}\cap B)\leq kd(w,\mathcal{M}\cap B)\qquad\text{ and }\qquad\phi(w,t)\in U,$
where $d(w,\mathcal{M}\cap B)=\inf_{x\in\mathcal{M}\cap B}\|w-x\|_{2}$. [^4] The definition of falconer1983differentiation is stated differently but is implied by our notion of geometric stability, when taking $f(w)=\phi(w,t)$.
Let $x\in\mathcal{M}\cap B$. The Morse-Bott property (Assumption˜2) implies that there exists a neighbourhood $U(x)\subset B$ of $x$, such that $F$ satisfies the Polyak-Łojasiewicz (PL) inequality with constant $\frac{\eta}{2}$, thanks to the equivalences between both conditions [rebjock2024fast]
@@
In particular, there is some $\delta_{0}(x)>0$ such that $B(x,\delta_{0}(x))\subset U(x)$. Thanks to rebjock2024fast, we can even choose $\delta_{0}(x)$ small enough so that there are some $\alpha,\beta$ such that
@@
By boundedness of $B$, $\alpha$ and $\beta$ can be chosen independently of $x\in\mathcal{M}\cap B$ here.
Now let $w\in B(x,\delta_{0}(x))$ and define $T(w)=\inf\{t\geq 0\mid\phi(w,t)\not\in B(x,\delta_{0}(x))\}$. Necessarily, $T(w)>0$ and for any $t\in[0,T(w))$, Equation˜17 applies to $\phi(w,t)$, so that for any $t\in[0,T(w))$
@@
Moreover for any $t\in[0,T(w))$, Equation˜18 also applies, so that
@@
Also, note that $\|\phi(w,t)-x\|\leq\frac{\delta_{0}(x)}{2}$ and $B(x,\delta_{0}(x))\subset B$ implies that $d(\phi(w,t),\mathcal{M}\cap B)=d(\phi(w,t),\mathcal{M})$. From there, Equation˜19 implies for any $t\geq 0$ and $w\in B(x,\delta(x))$:
@@
In particular, for any $k\geq 0$, we can choose a sufficiently large $t$ such that $d(\phi(w,t),\mathcal{M}\cap B)\leq kd(w,\mathcal{M}\cap B)$.
By compactness of $\mathcal{M}\cap B$, there is a finite family of $(x_{i})_{i\in[K]}\in\mathcal{M}\cap B$ such that $\bigcup_{i\in[K]}B(x_{i},\frac{1}{2}\delta(x_{i}))\supseteq\mathcal{M}\cap B$. We then define $U(B)=\bigcup_{i\in[K]}B(x_{i},\delta(x_{i}))$, which is also a finite covering of $\mathcal{M}\cap B$. We then take $t$ large enough such that for any $w\in U(B)$, $d(\phi(w,t),\mathcal{M}\cap B)\leq kd(w,\mathcal{M}\cap B)$ for $k<\frac{\min_{i\in[K]}\delta(x_{i})}{\max_{i\in[K]}\delta(x_{i})}$. In particular, our choice of $k$ is such that, for $f:w\mapsto\phi(w,t)$, $U(B)$ is invariant by $f$. Indeed, note that for any $w\in U(B)$,
$d(f(w),\mathcal{M}\cap B)$ $\leq kd(w,\mathcal{M}\cap B)$
$\leq\frac{\min_{i\in[K]}\delta(x_{i})}{\max_{i\in[K]}\delta(x_{i})}d(w,\mathcal{M}\cap B)$
$\leq\frac{1}{2}\min_{i\in[K]}\delta(x_{i}).$
In other words, there is $x\in\mathcal{M}\cap B$ such that $\|f(w)-x\|\leq\frac{1}{2}\min_{i\in[K]}\delta(x_{i})$. Moreover since $\bigcup_{i\in[K]}B(x_{i},\frac{1}{2}\delta(x_{i}))$ is a covering of $\mathcal{M}\cap B$, there is $j\in[K]$ such that $\|x-x_{j}\|<\frac{1}{2}\min_{i\in[K]}\delta(x_{i})$ and by triangle inequality:
$\|w-x_{j}\|<\delta(x_{j}),$
i.e., $w\in U(B)$.
Since $U(B)$ is invariant by $f$ and $k<1$, $\mathcal{M}\cap B$ is geometrically stable, so that we can apply falconer1983differentiation. It then implies that $\Phi$ is $\mathcal{C}^{2}$ on $U(B)$. Taking an increasing sequence of open bounded sets $B_{n}$ covering whole $\mathbb{R}^{d}$, we can then define $U=\bigcup_{n}U(B_{n})$ and conclude that $\Phi$ is $\mathcal{C}^{2}$ on $U$.
@@ B.2 Minimality of $F$ on neighbourhood
Consider Assumptions˜1 and 2. Let $U$ be an open neighbourhood of $\mathcal{M}$ such that $\Phi$ is continuous on $U$, then necessarily for any $w\in U\setminus\mathcal{M}$, $F(w)=\sup_{x\in\mathcal{M}}F(x)$.
@@ B.3 Alternative equation for $\tilde{w}^{\circ}$
If Assumption˜2 holds, the unique solution of Equation˜6 also corresponds to the unique solution of the following equation:
@@
The first one is a consequence of the definition of the Riemannian gradient and the fact that $T_{\mathcal{M}}(w)=\mathrm{Ker}(\nabla^{2}F(w))$ [see e.g., boumal2023introduction, Theorem 3.15 with $\nabla F$ being the local defining function of $\mathcal{M}$]. The second one is given by li2021happens. ∎